When 60.0 mL of a 0.400 M solution of HNO3 (aq) is combined with 60.0 mL of a 0.400 M solution of NaOH(aq) in a coffee-cup calorimeter, the final temperature of the solution is measured to be 26.6 ˚C. The initial temperature of the solutions is 24.0 ˚C. Calculate qrxn in joules, assuming the specific heat capacity of the final solution is 3.90 J·g–1·˚C–1, the density of the final solution is 1.04 g/mL, and the calorimeter constant is 45 J/oC. (Hint: start with qrxn + qsoln + qcal = 0.)

The balanced chemical equation for this reaction is:

HNO3 (aq) + NaOH(aq) → H2O(l) + NaNO3 (aq)
Answer #1

Q(rxn) = - ( Q(sol) + Q (cal))

Total volume = 60.0 + 60.0 = 120.0 ml

Mass = volume * density

= 120.0 * 1.04

= 124.8 grams

Q ( sol ) = mass * specific heat * ∆T

= 124.8 * 3.90 * ( 26.6 - 24.0 )

= 1265.472 J

Q ( cal ) = C(cal) * ∆T

= 45 * ( 26.6 - 24.0)

= 117 J

Q total = 1265.472 + 117

= 1382.472 J

Qrxn = - 1382.472 J

If you have any questions please comment

If you satisfied with the solution please rate it thanks